The position of a squirrel running in a park is given by At t = 5.00 s, how far is the squirrel from its initial position?
Answer #1

The equation of squirrel's displacement is missing

Let's assume

Your given equation is

r(t) = (0.28*t + 0.036*t^2) i + (0.019*t^3) j

then at t = 0,

squirrel's position will be

r(0) = (0.28*0 + 0.036*0^2) i + (0.019*0^3) j

r(0) = 0 i + 0 j

at t = 5 sec

squirrel's position will be

r(5) = (0.28*5 + 0.036*5^2) i + (0.019*5^3) j

r(0) = 2.3 i + 2.375 j

total distance from initial position will be

|r| = sqrt [(2.3 - 0)^2 + (2.375 - 0)^2]

|r| = 3.306 m